# AttenDence: Maximizing Attention Confidence for Test Time Adaptation

### v1

@@ 1 Introduction
Recent TTA methods have primarily focused on entropy minimization over the output distribution, which encourages the model to make confident predictions at test time. While effective, this approach treats the feature extractor as a black box and ignores rich internal representations that could guide adaptation. Vision transformers (dosovitskiy2021an), which have become the dominant architecture for visual recognition, offer a particularly promising internal signal: attention mechanisms that explicitly capture spatial relationships and feature importance.
@@
First, entropy minimization on output predictions works because confident predictions tend to be correct predictions. We hypothesize that this principle extends to attention mechanisms: confident attention patterns (low entropy) should correspond to better feature representations.
@@ 1.1 Contributions
  • We demonstrate that this approach substantially improves robustness on CIFAR-10-C while not hurting on clean data and slightly improving on it.
@@ 2.1 Test-Time Adaptation
Figure 3: Adam with the moments carried over between samples seems to help accuracy.
Figure 4: With regular SGD updates and resetting, there is an improvement in accuracy.
@@ 3.1 Problem Setting
We consider the standard test-time adaptation setting. Given a model $f_{\theta}$ trained on a source distribution $\mathcal{D}_{source}$, we encounter a sample $x_{i}$ from a shifted distribution $\mathcal{D}_{shift}$ at test time.
@@
    Generate prediction $\hat{y}_{i}=f_{\theta}(x_{i})$ using the adapted model
@@
We focus on the fully test-time setting where no source data, validation data, or labels are available during adaptation. This is the most challenging and practical setting for TTA.
@@ 3.2.1 Mathematical Formulation
Let $\mathbf{A}\in\mathbb{R}^{B\times H\times T\times T}$ denote the attention tensor from the final transformer layer, where $B$ is the batch size (1 here), $H$ is the number of attention heads, $T$ is the sequence length (CLS token + register tokens + patch tokens)
Let $t_{\text{cls}}$ denote the index of the CLS token, and let $t_{r}$ be the number of register or special tokens (including the CLS token itself). The post-softmax attention scores from the CLS token to patch tokens (i.e., tokens beyond the register tokens) is extracted as:
$\mathbf{A}_{\text{cls}}=\left[\mathbf{A}_{b,h,t_{\text{cls}},j}\right]_{\begin{subarray}{c}h=1,\dots,H\\ j=t_{r}+1,\dots,T\end{subarray}}\in\mathbb{R}^{H\times(T-t_{r})}$
We average across attention heads:
$\bar{\mathbf{a}}=\frac{1}{H}\sum_{h=1}^{H}\mathbf{A}_{\text{cls}}[h,:]$
Normalize to form a probability distribution since we exclude register tokens:
$\tilde{\mathbf{a}}=\frac{\bar{\mathbf{a}}}{\sum_{i=1}^{T-t_{r}}\bar{\mathbf{a}}_{i}}$
The attention entropy loss is then defined as:
$\mathcal{L}_{\text{attn}}=-\sum_{i=1}^{T-t_{r}}\tilde{\mathbf{a}}_{i}\log(\tilde{\mathbf{a}}_{i})$
Minimizing this loss encourages the attention from the CLS token to be focused (i.e., low-entropy) over patch tokens. We do episodic adaptation where the parameters are reset after each sample. For optimizer state we explore both episodic and continual cases.
3.2.2 Intuition and Design Choices
The key intuition behind attention entropy minimization is that confident attention patterns correspond to better feature representations. When a vision transformer attends strongly to the objects of interest, it can extract more discriminative features than when attention is spread uniformly across all patches.
@@ 3.3 Optimization and Adaptation Protocol
We explore Adam (kingma2015adam) and SGD to perform the single update. Interestingly, we observed that not resetting Adam’s optimizer state (first and second order moments) between samples can improve performance, even though we reset the model parameters. This could be due to the moments smoothing and scaling gradients based on their historical variance provides a more stable gradient during update.
@@ 4.1 Experimental Setup
We evaluate on severity level 5 (the most severe) for all corruption types. This provides a challenging test of robustness, as the corruptions are strong enough to significantly degrade performance.
@@
Evaluation Protocol: For each corruption type, we use 1,000 samples for hyperparameter search and then evaluate on 9,000 test. We report per-corruption accuracy and mean corruption accuracy (mCA) across all corruption types. Each sample (batch size 1) is adapted independently with a single gradient step.
Baselines: We evaluate without any test-time updates, providing a measure of the model’s inherent robustness, and perform TTA on clean data to check if it hurts accuracy.
@@ 4.2 Main Results
We hypothesize that momentum accumulation acts as temporal smoothing across the sequence of test samples. This finding is somewhat surprising given that we reset model parameters between samples. The optimizer momentum remembers gradients from previous samples even though the parameters themselves are reset. This creates a hybrid between episodic and continual adaptation: predictions are independent (since parameters reset), but the optimization trajectory is influenced by previous samples.
@@ 4.3.1 Architecture Dependence
4.3.2 Attention Quality Dependency
The effectiveness of attention entropy minimization depends on the quality of pretrained attention maps. We chose DINOv3 specifically because its self-supervised pretraining produces rich, interpretable attention patterns. Transformers with less interpretable attention (e.g., those trained purely on supervised ImageNet classification without registers) may benefit less from attention-based adaptation.

### v2

@@ 1 Introduction
Recent TTA methods have focused on entropy minimization over the output distribution, which encourages the model to make confident predictions at test time. While effective, this approach treats the feature extractor as a black box and ignores rich internal representations that could guide adaptation. Vision transformers (dosovitskiy2021imageworth16x16words), which have become the dominant architecture for visual recognition due to their scalability, offer a particularly promising internal signal: attention mechanisms that explicitly capture spatial relationships and feature importance.
@@
First, entropy minimization on output predictions works because confident predictions tend to be correct predictions. We hypothesize that this principle extends to attention mechanisms.
@@ 1.1 Contributions
  • We demonstrate that this approach substantially improves robustness on CIFAR-10-C while not hurting on clean data.
@@ 2.1 Test-Time Adaptation
Figure 3: Adam with the moments carried over between samples seems to help accuracy. Streaming (Batch size 1). Full results in Appendix B.
Figure 4: With regular SGD updates and resetting, there is an improvement in accuracy. Streaming (Batch size 1). Full results in Appendix B.
@@ 3.1 Problem Setting
We consider an episodic test-time adaptation setting. Given a model $f_{\theta}$ trained on a source distribution $\mathcal{D}_{source}$, we encounter a sample $x_{i}$ from a shifted distribution $\mathcal{D}_{shift}$ at test time.
@@
    Generate prediction $\hat{y}_{i}=f_{\theta^{\prime}}(x_{i})$ using the adapted model
@@ 3.2.1 Mathematical Formulation
Attention Tensor.
Let $\mathbf{A}(x_{i})\in\mathbb{R}^{H\times T\times T}$ be the post-softmax attention tensor from the final transformer layer and:
  • $x_{i}$ is the input sample.
  • $H$ is the number of attention heads,
  • $T$ is the sequence length (CLS + register + patch tokens),
  • $t_{\text{cls}}$ is the index of the CLS token,
  • $t_{r}$ is the number of special tokens (like register tokens), including the CLS token.
CLS-to-Patch Attention Extraction.
We extract attention scores from the CLS token to the patch tokens:
$\mathbf{A}_{\text{cls}}(x_{i})=\left[\mathbf{A}_{h,t_{\text{cls}},j}(x_{i})\right]_{\begin{subarray}{c}h=1,\dots,H\\ j=t_{r}+1,\dots,T\end{subarray}}\in\mathbb{R}^{H\times P}$
where $P=T-t_{r}$.
We then normalize to form a valid probability distribution after removing special tokens.
$a^{(h)}_{j}=\frac{\mathbf{A}_{h,t_{\text{cls}},j}(x_{i})}{\sum_{k=t_{r}+1}^{T}\mathbf{A}_{h,t_{\text{cls}},k}(x_{i})}\quad\text{for }j=t_{r}+1,\dots,T$
We then calculate the entropy per head.
$\mathcal{H}^{(h)}(x_{i})=-\sum_{j=1}^{P}a^{(h)}_{j}(x_{i})\log a^{(h)}_{j}(x_{i})$
The final loss is the average of the entropies per head:
$\mathcal{L}(x_{i})=\frac{1}{H}\sum_{h=1}^{H}\mathcal{H}^{(h)}(x_{i})$
$\frac{1}{H}\sum_{h=1}^{H}\sum_{j=1}^{P}a^{(h)}_{j}(x_{i})\log a^{(h)}_{j}(x_{i})$
Minimizing this loss encourages each attention head to place concentrated (low-entropy) focus on a smaller subset of patch tokens, rather than distributing attention diffusely. Averaging the distributions first, then taking their entropy, was tried but performed worse. This is reasonable as heads tend to specialize (raghu2022visiontransformerslikeconvolutional).
3.2.2 Design Choices
@@ 3.3 Optimization and Adaptation Protocol
We explore Adam (kingma2015adam) and Stochastic gradient Descent (SGD) to perform the single update. Interestingly, we observed that not resetting Adam’s optimizer state (first and second order moments) between samples can improve performance, even though we reset the model parameters. We discuss this further in Appendix A.
@@ 4.1 Experimental Setup
Evaluation Protocol: For each corruption type, we use 1,000 samples for hyperparameter search and then evaluate on 9,000 test samples. We report per-corruption accuracy and mean corruption accuracy (mCA) across all corruption types. Each sample (batch size 1) is adapted independently with a single gradient step.
Baselines: We evaluate without any test-time updates, providing a measure of the model’s inherent robustness, and perform TTA on clean data to show that this method does not hurt on clean data.
@@ 4.2 Main Results
We hypothesize that Adam’s moments act as temporal smoothing across the sequence of test samples. This creates a hybrid between episodic and continual adaptation: parameters are reset, but previous samples influence the optimization trajectory. We observed that the distribution of the attention scores from CLS to image patches does have a tail across all heads. Our hypothesis is that this helps Adam adapt, or rather hurts SGD, as shown by kunstner2024heavytailedclassimbalanceadam. This is discussed further in Appendix A
@@ Acknowledgements
Appendix A Appendix - Optimizer Moments
We saw that resetting the model weights per sample after adaptation but not resetting the optimizer moments/state across samples helps significantly.
We observed that the distribution of the attention scores from CLS to image patches does have a tail across all heads. Our hypothesis is that this is what Adam adapts to (kunstner2024heavytailedclassimbalanceadam) and may explain the performance gap between regular SGD and Adam with only moments carried over. While the log-log plots do not show a clear power law distribution, they exhibit characteristics consistent with heavy-tailed behavior.
Figure 5: Log-log plot of attention weights vs frequency over 10% of the data.
Appendix B Appendix - Full Results & Hyperparameters
B.1 Adam with Moments Carried Over
Hyperparameters
After using 10% of the data for hyperparameter search, a learning rate of $1\times 10^{-4}$ was selected, and Adam was configured with exponential moving–average decay rates $\beta_{1}=0.99$ and $\beta_{2}=0.999$. The corruptions were visited in alphabetical order of corruption name.
Table 1: Per-corruption accuracy on 90% of CIFAR 10-C (%) — Adam (Moments Carried Over) — Streaming (Batch size 1).
Corruption Brightness Contrast Defocus Blur Elastic Transform Fog Frost Gaussian Blur Gaussian Noise
Accuracy (%) 95.71 93.79 89.11 72.90 79.86 89.54 89.39 71.40
Corruption Glass Blur Impulse Noise JPEG Compression Motion Blur Pixelate Saturate Shot Noise Snow
Accuracy (%) 61.57 86.12 81.32 84.80 82.10 95.19 74.90 90.12
Corruption Spatter Speckle Noise Zoom Blur Mean Accuracy
Accuracy (%) 95.37 75.96 90.59 84.20
B.2 SGD
After using 10% of the data for hyperparameter search, a learning rate of $1\times 10^{-3}$ was selected.
Table 2: Per-corruption accuracy on 90% of CIFAR 10-C (%) — SGD (Truly Episodic) — Streaming (Batch size 1)
Corruption Brightness Contrast Defocus Blur Elastic Transform Fog Frost Gaussian Blur Gaussian Noise
Accuracy (%) 94.72 89.32 83.09 64.03 66.34 88.03 82.48 63.10
Corruption Glass Blur Impulse Noise JPEG Compression Motion Blur Pixelate Saturate Shot Noise Snow
Accuracy (%) 53.67 82.27 77.61 77.02 77.79 93.83 65.18 87.39
Corruption Spatter Speckle Noise Zoom Blur Mean Accuracy
Accuracy (%) 94.52 67.16 84.62 78.54
B.3 Clean data
These results show that these approaches do not hurt accuracy on clean data and slightly improve it.
Table 3: Results on 90% of CIFAR-10 test set.
Method Accuracy (%)
Clean accuracy 96.66
Adam (moments carried over) 97.4
SGD (truly episodic) 97.16